If i have a 84.80 right now in my class and I don't do my final exam which is worth 10% and I get a zero what well it bring my grade down too?
Answer by rothauserc(4718)

0.90 * ((84+80)/2) = 73.8
:
0.10 * (0) = 0
:
****************************
your grade = 73.8 + 0 = 73.8
****************************
: